# A Non-Commutative Kalb-Ramond Black Hole

### v2

@@ IX.3 Time delay of light
$\left(\frac{\mathrm{d}r}{\mathrm{d}t}\right)^{2}=\frac{A^{(\Theta,\ell)}D^{(\Theta,\ell)}-\frac{\ell^{2}}{E^{2}}A^{2}(r)}{B^{(\Theta,\ell)}D^{(\Theta,\ell)}}\,.$ (134)
@@
High–precision data from the Cassini spacecraft Bertotti:2003rm ; Will:2014kxa have led to the most stringent bound on the PPN parameter $\gamma$, constraining it to $|\gamma-1|<2.3\times 10^{-5}$. In natural units, the average Earth–Sun separation is defined as one astronomical unit (AU), corresponding to $r_{E}=1\,\text{AU}=2.457\times 10^{45}$. During the measurement, the Cassini probe was positioned at $r_{R}=8.46\,\text{AU}$ from the Sun, while the light signal’s minimum radial distance from the Sun was $r_{\text{min}}=1.6,R_{\odot}$, with the solar radius given by $R_{\odot}=4.305\times 10^{43}$. Using these parameters, one obtains an upper bound on the non–commutative parameter: $|\Theta^{2}|\leq 5.001\times 10^{14},\text{m}^{2}$.
@@
Mercury precession $-2976.57\,\text{m}^{2}\leq\Theta^{2}\leq 595.315\,\text{m}^{2}$ $-1.817\times 10^{-11}\leq\ell\leq 3.634\times 10^{-12}$
Light deflection $-3.872\times 10^{13}\,\text{m}^{2}\leq\Theta^{2}\leq 1.936\times 10^{14}\,\text{m}^{2}$ $-1.333\times 10^{-5}\leq\ell\leq 6.667\times 10^{-5}$
Shapiro time delay $-5.001\times 10^{14}\,\text{m}^{2}\leq\Theta^{2}\leq 5.001\times 10^{14}\,\text{m}^{2}$ $-5.700\times 10^{-6}\leq\ell\leq 5.700\times 10^{-6}$
@@ X Conclusion
We also analyzed the thermodynamic behavior of the system. In contrast to the findings of Ref. Juric:2025kjl , where the surface gravity for the Schwarzschild solution under the same Moyal twist was shown to be ill–defined—therefore, obstructing a consistent thermodynamic treatment. On the other hand, in our Kalb–Ramond framework, the surface gravity was well–defined instead. As a result, the thermodynamic quantities could be properly evaluated. We calculated the Hawking temperature $T^{(\Theta,\ell)}$ and the heat capacity $C_{V}^{(\Theta,\ell)}$. Since the entropy remained $\pi r_{h}^{2}$ and the horizon was unaffected by $\Theta$, an explicit discussion of the entropy was omitted.
@@
Finally, the constraints obtained from solar system observations establish the following bounds: from Mercury’s perihelion precession, we got $-2976.57,\text{m}^{2}\leq\Theta^{2}\leq 595.315\,\text{m}^{2}$ and $-1.817\times 10^{-11}\leq\ell\leq 3.634\times 10^{-12}$; from light deflection, the allowed range was $-3.872\times 10^{13}\,\text{m}^{2}\leq\Theta^{2}\leq 1.936\times 10^{14}\,\text{m}^{2}$ and $-1.333\times 10^{-5}\leq\ell\leq 6.667\times 10^{-5}$; and from the Shapiro time delay, we had $-5.001\times 10^{14}\,\text{m}^{2}\leq\Theta^{2}\leq 5.001\times 10^{14}\,\text{m}^{2}$ and $-5.700\times 10^{-6}\leq\ell\leq 5.700\times 10^{-6}$.

### v3

@@ IX.3 Time delay of light
$\left(\frac{\mathrm{d}r}{\mathrm{d}t}\right)^{2}=\frac{A^{(\Theta,\ell)}D^{(\Theta,\ell)}-\frac{\ell^{2}}{E^{2}}{(A^{(\Theta,\ell)})}^{2}}{B^{(\Theta,\ell)}D^{(\Theta,\ell)}}\,.$ (137)
@@
High–precision data from the Cassini spacecraft Bertotti:2003rm ; Will:2014kxa have led to the most stringent bound on the PPN parameter $\gamma$, constraining it to $|\gamma-1|<2.3\times 10^{-5}$. In natural units, the average Earth–Sun separation is defined as one astronomical unit (AU), corresponding to $r_{E}=1\,\text{AU}=2.457\times 10^{45}$. During the measurement, the Cassini probe was positioned at $r_{R}=8.46\,\text{AU}$ from the Sun, while the light signal’s minimum radial distance from the Sun was $r_{\text{min}}=1.6,R_{\odot}$, with the solar radius given by $R_{\odot}=4.305\times 10^{43}$. Using these parameters, one obtains an upper bound on the non–commutative parameter: $|\Theta^{2}|\leq 5.001\times 10^{14}\,\text{m}^{2}$.
@@
Mercury precession
$2976.57\,\text{m}^{2}\leq\Theta^{2}\leq 595.315\,\text{m}^{2}$
$1.817\times 10^{-11}\leq\ell\leq 3.634\times 10^{-12}$
Light deflection
$3.872\times 10^{13}\,\text{m}^{2}\leq\Theta^{2}\leq 1.936\times 10^{14}\,\text{m}^{2}$
$1.333\times 10^{-5}\leq\ell\leq 6.667\times 10^{-5}$
Shapiro time delay
$5.001\times 10^{14}\,\text{m}^{2}\leq\Theta^{2}\leq 5.001\times 10^{14}\,\text{m}^{2}$
$5.700\times 10^{-6}\leq\ell\leq 5.700\times 10^{-6}$
@@ X Conclusion
We also examined the thermodynamic properties of the system. Unlike the result reported in Ref. Juric:2025kjl , where the Schwarzschild case under the same Moyal twist led to an ill–defined surface gravity—preventing a consistent thermodynamic analysis—our Kalb–Ramond framework yielded a well–defined surface gravity. Consequently, the thermodynamic quantities could be consistently determined. In particular, we computed the Hawking temperature $T^{(\Theta,\ell)}$ and the heat capacity $C_{V}^{(\Theta,\ell)}$. Since the entropy remained $\pi r_{h}^{2}$ and the horizon was not modified by $\Theta$, a separate discussion of the entropy was not included.
@@
Finally, the constraints obtained from solar system observations establish the following bounds: from Mercury’s perihelion precession, we got $2976.57\,\text{m}^{2}\leq\Theta^{2}\leq 595.315\,\text{m}^{2}$ and $1.817\times 10^{-11}\leq\ell\leq 3.634\times 10^{-12}$; from light deflection, the allowed range was $3.872\times 10^{13}\,\text{m}^{2}\leq\Theta^{2}\leq 1.936\times 10^{14}\,\text{m}^{2}$ and $1.333\times 10^{-5}\leq\ell\leq 6.667\times 10^{-5}$; and from the Shapiro time delay, we had $5.001\times 10^{14}\,\text{m}^{2}\leq\Theta^{2}\leq 5.001\times 10^{14}\,\text{m}^{2}$ and $5.700\times 10^{-6}\leq\ell\leq 5.700\times 10^{-6}$.